You have been offered a project paying $300 at the beginning of each year for the next 20 years. What is the maximum amount of money you would invest in this project if you expect 9 percent rate of return to your investment?
Answer #1

In this question, I have to calculate the present value of the amounts which I will be receiving in future. I am receiving $300 at the beginning of each year for 20 years.

Present value= PV of $300 at starting of year 1 + PV $300 every year for 19 years

Present value= 300*1 + 300* PVAF(9%,19)

Present value= 300+300*8.9501

Present value = 300+2685.03= $2985.03

Thus, the maximum amount that I would invest in this project is $2985.03 since at this amount my NPV is equal to zero. The value of my inflows and ouflows will be equal.

PVAF value has been taken since we are receiving the same amount. PVAF @ 9% for 19 years is 9.8501